OpenStudy (anonymous):

Find thw anti-derivative of \[f(x)=\frac{ -1 }{ x^2 }e ^{1/x}\]

OpenStudy (johnweldon1993):

\[\large -\int \frac{e^{1/x}}{x^2}\] Make a u-sub of 1/x

OpenStudy (johnweldon1993):

Oh...what have you learned?

OpenStudy (anonymous):

How do you write in this format? f(x)=F(x)+C

zepdrix (zepdrix):

\[\Large\rm \int\limits f(x)dx=F(x)+C\]Where F(x) is the anti-derivative of f(x). Did you understand the u-substitution?

OpenStudy (anonymous):

Is there another way to solve it without using the u-substitution?

OpenStudy (freckles):

i can't think of a better way

OpenStudy (anonymous):

The reason that I ask that was because the question was at the beginning of antiderivatives/integration. I learned the u-substitution as a last topic...

ganeshie8 (ganeshie8):

you may use advanced guessing what is the derivative of \(\large e^{1/x}\) ?

OpenStudy (anonymous):

e^1/x

ganeshie8 (ganeshie8):

what do you know about chain rule

OpenStudy (anonymous):

Chain rule applies to composite function. It describe of outer evulated at inner, times derivative of inner

OpenStudy (johnweldon1993):

So using that logic...if we consider the "inner" as 1/x The derivative of \(\large e^{1/x}\) would be \(\large \frac{d(1/x)}{dx} \times e^{1/x}\) So what is the derivative of \(\large 1/x\) ?

OpenStudy (anonymous):

lnx?

OpenStudy (freckles):

you found the antiderivative for positive x

OpenStudy (freckles):

that you found the antiderivative of 1/x for positive x the question was to differentiate

ganeshie8 (ganeshie8):

derivative if lnx is 1/x derivative of 1/x is not lnx

OpenStudy (anonymous):

ooh... derivative of 1/x is -1/x^2 ...

OpenStudy (johnweldon1993):

Correct...and to finish up from the previous post \[\large \frac{d}{dx}e^{1/x} = e^{1/x} \times \frac{d}{dx}\frac{1}{x}\] Since we just found the later \(\large \frac{d}{dx} \frac{1}{x} = -\frac{1}{x^2}\) we have \[\large \frac{d}{dx}e^{1/x} = -\frac{e^{1/x}}{x^2}\] right? Now what do you notice?

OpenStudy (anonymous):

So F(x)= e^{1/x} +C ?

OpenStudy (johnweldon1993):

Perfect! :)

OpenStudy (anonymous):

Thank you! :) I got it now. I was messed up b/w derivatives and anti-derivatives. My bad...

OpenStudy (johnweldon1993):

Yeah sometimes they can be hard to keep straight...but as long as you can derive them you're all set!